OpenStudy (pooh):

Soybean meal is 16% protein; Cornmeal is 8% protein. How many pound of each should be mixed together to get a 320lb that is 13% protein

OpenStudy (anonymous):

i think i have seen this so many times i could do it in my sleep. your equation is \[.16x+.08(320-x)=320\times .13\]

OpenStudy (anonymous):

because if x represents soybeans then it contributes .16 x amount of protein. and if x is amount of soybeans and total is 320 that leaves 320-x of cornmeal which has .08(320-x) protein. the total is 13% of 320 which is .13 times 320 now to solve

OpenStudy (pooh):

so x is 200

OpenStudy (anonymous):

get \[.16x+25.6-.08x=41.6\] yes you got it!

OpenStudy (pooh):

but that only gives me one answer I need to know how much soybean and how much corn meal

OpenStudy (anonymous):

Soybean 200 Cornmeal 120

OpenStudy (anonymous):

i put x = amount of soybeans and found x = 200 since the total is 320 we know the rest yest?

OpenStudy (anonymous):

I solve this problem a few minute ago, let me find and copy for you

OpenStudy (pooh):

Thanks both of you have been a great help

OpenStudy (anonymous):

urwc